OpenStudy (anonymous):

Please factor n^3 - 3n^2 -168n + 720 = 0

OpenStudy (anonymous):

\[n^3 - 3n^2 - 168n + 720 = 0\]

OpenStudy (anonymous):

And show me how you did it too.

OpenStudy (campbell_st):

I used trial and error for find f(12) = 0 then (n -12) is a factor then use polynomial division \[(n^3 - 3n^2 - 168n + 720)/(x -12) = (n - 12)(n^2 + 9n - 60)\] to get the other factor